OpenStudy (anonymous):

A mass of 2kg, initially at rest, slides down a frictionless inclined plane of length 1.5 meters in 2.75 seconds. What was the angle of inclination of the plane?

OpenStudy (anonymous):

It can be anything since you said it is frictionless

OpenStudy (anonymous):

Well, I'm not good at explanation :( I hope you get it.

OpenStudy (anonymous):

we dont know the angle of inclination, but we do know the distance and time taken to cover that distance which we can use to find the average speed: velocity = 1.5/2.75 = 0.5454 m/s find the acceleration vf=vi+at, since initial velocity was zero, vi = 0, thereforce: vf=at or a = vf/t = 0.5454/2.75 = 0.1983 m/s^2 \[mgsin\theta=a\] you already have acceleration \[sin\theta=a/mg\] \[\theta=sin^{-1}(a/mg)\] \[\theta= 1.59^{o}\] does it look logical ??, how can 2 kg mass slide at that speed/acceleration on 1.59 degree of inclination ?? (maybe correct becoz we assumed its friction less, maybe my calculations are wrong)

OpenStudy (anonymous):

oops, i used 0.5454 as acceleration in my calculation, using 0.1983 m/s^2 i get even smaller angle 0.579 degree

OpenStudy (anonymous):

16.5 degree